What are stationary waves? Discuss the formation of stationary waves at a free boundary, and hence explain their changes with respect to the position and time. 


1
Expert's answer
2021-02-11T08:54:54-0500

The stationary wave, is a wave which oscillates in time but whose peak amplitude profile does not move in space. The equation of standing wave on an infinite string is:


"y(x,t) = A\sin\left( \dfrac{2\pi x}{\lambda} \right)\cos(\omega t)"

where "A" is an amplitude, "\lambda" is a wavelength, "\omega" is an angular frequency, and "x,t" are position and time coordinates respectively. Considering the string with one fixed "x = 0" and one free "x = L" ends. Thus, at the fixed end we have the node "y(0,t) = 0", at the free end we have tha anti-node "y = y_{max}". Thus, it should be:


"\sin\left( \dfrac{2\pi L}{\lambda} \right) = 1"

Expressing the wavelength, obtain:


"\lambda= \dfrac{4L}{n}, \space \space n = 1,3,5..."

Answer. Thus, we have antinode at the free end, that oscillates in time as "\cos(\omega t)".